# Asymptotics for conformal inference

### v1

@@ Appendix D Weighted novelty detection for a general weight function
$\sqrt{\tau_{n,m}}\left(\mathrm{TDP}_{m}^{(n)}\left({\mathcal{R}}^{w}_{\alpha}\right)-G^{w}(\mathcal{T}^{w}_{\alpha})\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{({G^{w}})^{\prime}(\mathcal{T}^{w}_{\alpha})}{\frac{1}{\alpha}-({G_{{\tiny\mbox{mixt}}}^{w}})^{\prime}(\mathcal{T}^{w}_{\alpha})}\mathbb{Z}^{w}({\mathcal{T}^{w}_{\alpha}})+\mathbb{Z}^{w}_{1}(\mathcal{T}^{w}_{\alpha}),$
@@ E.1 Donsker and Glivenko-Cantelli theorems
The results below can be found in van der Vaart, (1998) and Shorack and Wellner, (1986).
@@
[…]
@@
Let $(U_{k})_{k\geqslant 1}$ be i.i.d. random variables uniformly distributed over $(0,1)$. Denote $\mathbb{F}^{(n)}$ the e.c.d.f. the family $\left(U_{k}\right)_{k\in\llbracket n\rrbracket}$. Then,
[…]
Theorem E.3 (Donsker, (1952)).
Let $(X_{i})_{i\geqslant 1}$ iid real random variables. Let $F$ the cumulative distribution function of $X_{1}$ and $\mathbb{F}_{n}$ the empirical cdf of $\{X_{1},\cdots,X_{n}\}$. Then,
$\sqrt{n}\left(\mathbb{F}_{n}-F\right)\overset{\mathcal{L}}{\rightarrow}{\mathbb{U}_{F}},$
@@
Let $w$ a weight function satisfying Assumption 4. Then the family of function =$\mathcal{F}=\left\{w_{t}:x\in\mathbb{R}\mapsto w(x){\mathds{1}_{x\leqslant t}};t\in\mathbb{R}\right\}$ is $P_{\tiny\mbox{cal}}$-Donsker and $P_{\tiny\mbox{cal}}$-Glivenko-Cantelli.
@@ E.2 Hadamard differentiability
The three first results below can be found in van der Vaart, (1998).
@@
Let $F:[a,b]\rightarrow\mathbb{R}$ be a continuously differentiable function. The map $T:\phi\mapsto F\circ\phi$ with entries being functions $\phi:T\mapsto[a,b]$ contained in $\ell^{\infty}(T)$ is Hadamard differentiable tangentially to $\mathcal{C}(T)$ with derivative:
$\dot{T}_{\phi}(H)=F^{\prime}(\phi)H$
@@
The cumulative distribution function $F$ have a compact support $[a,b]$ in the sense that that for all $x\leqslant a$, $F(x)=0$, for all $x\geqslant b$, $F(x)=1$ and for all $x\in(a,b)$, $10<F(x)<$ 1. Furthermore, $F$ is continuously differentiable on $(a,b)$ with stricly positive derivative.
@@
Let $(X_{i})_{i\geqslant 1}$ be iid real random variables. Let $F$ be the cumulative distribution function of $X_{1}$ and $\mathbb{F}_{n}$ be the empirical cdf of $\{X_{1},\cdots,X_{n}\}$. Denote $F^{-1}$ and $\mathbb{F}^{-1}_{n}$ the corresponding quantile functions. Assume that $F$ satisfies Assumption 8 with derivative $f$. Then,
$\sqrt{n}\left(\mathbb{F}^{-1}_{n}-F^{-1}\right)\overset{\mathcal{L}}{\rightarrow}\frac{\mathbb{U}}{f\circ F^{-1}}=({F^{-1}})^{\prime}\mathbb{U},$
@@
$T(K+tH_{t})-T(K)$ $=\frac{K(1)\left(K+tH_{t}\right)-\left(K(1)+tH_{t}(1)\right)K}{\left(K(1)+tH_{t}(1)\right)K(1)}$
$=\frac{K(1)tH_{t}-tH_{t}(1)K}{\left(K(1)+tH_{t}(1)\right)K(1)}$
$=t\times\left(\frac{H_{t}}{K(1)+tH_{t}(1)}-\frac{H_{t}(1)}{K(1)+tH_{t}(1)}\frac{K}{K(1)}\right).$
@@
We gather below the formulas of the Hadamard derivatives of the functionals of interest when studying the asymptotic of the $\mathrm{FDP}$. They are obtained from Neuvial, (2008).
@@
$(\dot{\mathcal{T}}^{\mathrm{BH}_{\alpha}})_{G}(H)=\frac{H\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)\right)}{\alpha^{-1}-G^{\prime}\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)\right)}.$
@@
$\Phi$ $:(F_{0},F_{1})\rightarrow\pi F_{1}(\mathcal{T}^{\mathrm{BH}_{\alpha}}(F)),$ (48)
@@
$\dot{\nu}_{G_{0},G_{1}}(H_{0},H_{1})$ $=G_{0}^{\prime}\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)\right)\frac{H\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)\right)}{\alpha^{-1}-G^{\prime}\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)\right)}+H_{0}\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)\right);$
$\dot{\Psi}_{G_{0},G_{1}}(H_{0},H_{1})$ $=G_{1}^{\prime}\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)\right)\frac{H\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)\right)}{\alpha^{-1}-G^{\prime}\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)\right)}+H_{1}\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)\right);$
$\dot{\Phi}_{G_{0},G_{1}}(H_{0},H_{1})$ $=\frac{1}{G\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)\right)}\left(\dot{\nu}_{G_{0},G_{1}}(H_{0},H_{1})-\frac{\pi G_{0}\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)\right)}{G\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)\right)}(\dot{\mathcal{T}}^{\mathrm{BH}_{\alpha}})_{G}(H)\right),$
@@ E.3 Random change of time
We present here a version of the random change of time lemma of Billingsley, (1999) (page 151), which is adapted to the topological spaces $D(\mathbb{R})$, $\ell^{\infty}(0,1)$ and $D(0,1)$.
@@
Since $\mathbb{U}$ and $\mathbb{V}$ are continuous, $(U_{n},V_{n})_{n}$ converges in distribution to $(\mathbb{U},\mathbb{V})$ on ()×$\ell^{\infty}(\mathbb{R})\times\ell^{\infty}(0,1)$ by Lemma E.12. Hence, by Slutsky’s lemma, the sequence $(U_{n},V_{n},F_{n})_{n}$ converges in distribution to $(\mathbb{U},\mathbb{V},F)$ on ()×$\ell^{\infty}(\mathbb{R})\times(\ell^{\infty}(0,1))^{2}$. Hence by Lemma E.11 and the continuous mapping theorem we obtain,
@@
Let $U\in\mathcal{C}(\mathbb{R})$ and $(V,F)\in(\ell^{\infty}(0,1))^{2}$. Let $(U_{n},V_{n},F_{n})_{n}$ a sequence in ()×$\ell^{\infty}(\mathbb{R})\times(\ell^{\infty}(0,1))^{2}$ which converges to $(U,V,F)$. Let $K\subset(0,1)$ be a compact set. Let $\varepsilon>0$. The set $\left\{F_{n}(x),x\in K,n\in\mathbb{N}\right\}$ is bounded since $\left\|F_{n}-F\right\|_{\infty,K}\rightarrow 0$ and $(F_{n})_{n}$ and $F$ belongs in $\ell^{\infty}(0,1)$, hence is included in a compact set $K^{\prime}$ of $\mathbb{R}$. Since $U$ is continuous, we have $\eta>0$ such that for all $(x,y)\in{K^{\prime}}^{2}$, $\left|x-y\right|\leqslant\eta$ implies that $\left|U(x)-U(y)\right|\leqslant\varepsilon$. Let $N>0$ such that for all $n\geqslant N$, $\left\|U_{n}-U\right\|_{\infty,K^{\prime}}\leqslant\varepsilon$ and $\left\|F_{n}-F\right\|_{\infty,K}\leqslant\eta$. Then for all $n\geqslant N$,
$\left\|U_{n}\circ F_{n}-U\circ F\right\|_{\infty,K}$ $\leqslant\left\|U_{n}\circ F_{n}-U\circ F_{n}\right\|_{\infty,K}+\left\|U\circ F_{n}-U\circ F\right\|_{\infty,K}$
$\leqslant\left\|U_{n}-U\right\|_{\infty,K^{\prime}}+\varepsilon$
@@
where the first $\varepsilon$ appears since for all $x\in K$, $F_{n}(x)\in K^{\prime}$, $F(x)\in K^{\prime}$ and $\left|F_{n}(x)-F(x)\right|\leqslant\eta$. Furthermore for all $\tilde{K}$ a compact subset of $(0,1)$, $||V_{n}-V||_{\infty,\tilde{K}}\rightarrow 0$. Hence, $\Psi$ is continuous at $(U,V,F)$. ∎
Finally, the following result is classical and shows how a convergence in distribution on $D[0,1]$ can imply a uniform convergence (Billingsley,, 1999), with the local topology defined in Section 2.3.
Lemma E.12 (Billingsley, (1999)).
@@ Appendix F Useful Lemmas
Let $\left(s_{1},\cdots,s_{n+1}\right)\in{\mathbb{R}\cup\left\{+\infty\right\}}$, $(w_{k})_{k\in\llbracket n+1\rrbracket}$ be a family of positive weight summing to $1$ and $\mu=\sum_{k\in\llbracket n+1\rrbracket}w_{k}\delta_{s_{k}}$ be a probability measure on {$\mathbb{R}\cup\left\{+\infty\right\}$. Define for all $\alpha\in(0,1)$, $Q_{\alpha}(\mu)=\inf\left\{x\in\mathbb{R}:\mu\left([-\infty,x]\right)\geqslant\alpha\right\}$ the $\alpha$-quantile of the probability measure $\mu$. Then, for all $t\in\mathbb{R}$ and $\alpha\in(0,1)$, the two following assertions are equivalent:
@@
$S:=\max_{k\in\llbracket n+1\rrbracket}\left\{s_{k}:s_{k}<t\right\}\in\mathbb{R}$

### v3

@@ Appendix D Weighted novelty detection for a general weight function
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\mathrm{TDP}_{m}^{(n)}\mathopen{}\mathclose{{\left({\mathcal{R}}^{w}_{\alpha}}}\right)-G^{w}(\mathcal{T}^{w}_{\alpha})}}\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{({G^{w}})^{\prime}(\mathcal{T}^{w}_{\alpha})}{\frac{1}{\alpha}-({G_{{\tiny\mbox{mixt}}}^{w}})^{\prime}(\mathcal{T}^{w}_{\alpha})}\mathbb{Z}^{w}({\mathcal{T}^{w}_{\alpha}})+\mathbb{Z}^{w}_{1}(\mathcal{T}^{w}_{\alpha}),$
@@ E.1 Donsker and Glivenko-Cantelli theorems
The results below can be found in van der Vaart (1998) and Shorack and Wellner (1986).
@@
$\mathopen{}\mathclose{{\left\lVert{\widehat{F}}_{m,\tiny\mbox{test}}-F_{\tiny\mbox{test}}}}\right\rVert_{\infty}\rightarrow 0\text{ a.s.}$
@@
Let $(U_{k})_{k\geqslant 1}$ be i.i.d. random variables uniformly distributed over $(0,1)$. Denote $\mathbb{F}^{(n)}$ the e.c.d.f. the family $\mathopen{}\mathclose{{\left(U_{k}}}\right)_{k\in\llbracket n\rrbracket}$. Then,
$\mathopen{}\mathclose{{\left\lVert\mathopen{}\mathclose{{\left(\mathbb{F}^{(n)}}}\right)^{-1}-I}}\right\rVert_{\infty}\rightarrow 0\text{ a.s.}$
Theorem E.3 (Donsker (1952)).
Let $(X_{i})_{i\geqslant 1}$ i.i.d. real random variables. Let $F$ the cumulative distribution function of $X_{1}$ and $\mathbb{F}_{n}$ the empirical cdf of $\{X_{1},\cdots,X_{n}\}$. Then,
$\sqrt{n}\mathopen{}\mathclose{{\left(\mathbb{F}_{n}-F}}\right)\overset{\mathcal{L}}{\rightarrow}{\mathbb{U}_{F}},$
@@
Let $w$ a weight function satisfying Assumption 4. Then the family of function $\mathcal{F}=\mathopen{}\mathclose{{\left\{w_{t}:x\in\mathbb{R}\mapsto w(x){\mathds{1}_{x\leqslant t}};t\in\mathbb{R}}}\right\}$ is $P_{\tiny\mbox{cal}}$-Donsker and $P_{\tiny\mbox{cal}}$-Glivenko-Cantelli.
@@ E.2 Hadamard differentiability
The three first results below can be found in van der Vaart (1998).
@@
Let $F:[a,b]\rightarrow\mathbb{R}$ be a continuously differentiable function. The map $T:\phi\mapsto F\circ\phi$ with entries being functions $\phi:E\mapsto[a,b]$ contained in $\ell^{\infty}(E)$ is Hadamard differentiable tangentially to $\ell^{\infty}(E)$ with derivative:
$\dot{J}_{\phi}(H)=F^{\prime}(\phi)H.$
@@
The cumulative distribution function $F$ have a compact support $[a,b]$ in the sense that that for all $x\leqslant a$, $F(x)=0$, for all $x\geqslant b$, $F(x)=1$ and for all $x\in(a,b)$, $0<F(x)<1$. Furthermore, $F$ is continuously differentiable on $(a,b)$ with stricly positive derivative.
@@
Let $(X_{i})_{i\geqslant 1}$ be i.i.d. real random variables. Let $F$ be the cumulative distribution function of $X_{1}$ and $\mathbb{F}_{n}$ be the empirical cdf of $\{X_{1},\cdots,X_{n}\}$. Denote $F^{-1}$ and $\mathbb{F}^{-1}_{n}$ the corresponding quantile functions. Assume that $F$ satisfies Assumption 8 with derivative $f$. Then,
$\sqrt{n}\mathopen{}\mathclose{{\left(\mathbb{F}^{-1}_{n}-F^{-1}}}\right)\overset{\mathcal{L}}{\rightarrow}\frac{\mathbb{U}}{f\circ F^{-1}}=({F^{-1}})^{\prime}\mathbb{U},$
@@
$T(K+tH_{t})-T(K)$ $=\frac{K(1)\mathopen{}\mathclose{{\left(K+tH_{t}}}\right)-\mathopen{}\mathclose{{\left(K(1)+tH_{t}(1)}}\right)K}{\mathopen{}\mathclose{{\left(K(1)+tH_{t}(1)}}\right)K(1)}$
$=\frac{K(1)tH_{t}-tH_{t}(1)K}{\mathopen{}\mathclose{{\left(K(1)+tH_{t}(1)}}\right)K(1)}$
$=t\times\mathopen{}\mathclose{{\left(\frac{H_{t}}{K(1)+tH_{t}(1)}-\frac{H_{t}(1)}{K(1)+tH_{t}(1)}\frac{K}{K(1)}}}\right).$
@@
We gather below the formulas of the Hadamard derivatives of the functionals of interest when studying the asymptotic of the $\mathrm{FDP}$. They are obtained from Neuvial (2008).
@@
$(\dot{\mathcal{T}}^{\mathrm{BH}_{\alpha}})_{G}(H)=\frac{H\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)}}\right)}{\alpha^{-1}-G^{\prime}\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)}}\right)}.$
@@
$\Phi$ $:(F_{0},F_{1})\rightarrow F_{1}(\mathcal{T}^{\mathrm{BH}_{\alpha}}(F)),$ (54)
@@
$\dot{\nu}_{G_{0},G_{1}}(H_{0},H_{1})$ $=G_{0}^{\prime}\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)}}\right)\frac{H\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)}}\right)}{\alpha^{-1}-G^{\prime}\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)}}\right)}+H_{0}\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)}}\right);$
$\dot{\Psi}_{G_{0},G_{1}}(H_{0},H_{1})$ $=\frac{\pi}{G\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)}}\right)}\mathopen{}\mathclose{{\left(\dot{\nu}_{G_{0},G_{1}}(H_{0},H_{1})-\frac{G_{0}\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)}}\right)}{G\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)}}\right)}\mathopen{}\mathclose{{\left[H\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)}}\right)+G^{\prime}\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)}}\right)(\dot{\mathcal{T}}^{\mathrm{BH}_{\alpha}})_{G}(H)}}\right]}}\right);$
$\dot{\Phi}_{G_{0},G_{1}}(H_{0},H_{1})$ $=G_{1}^{\prime}\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)}}\right)\frac{H\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)}}\right)}{\alpha^{-1}-G^{\prime}\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)}}\right)}+H_{1}\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)}}\right),$
@@ E.3 Random change of time
We present here a version of the random change of time lemma of Billingsley (1999) (page 151), which is adapted to the topological spaces $D(\mathbb{R})$, $\ell^{\infty}(0,1)$ and $D(0,1)$.
@@
Since $\mathbb{U}$ and $\mathbb{V}$ are continuous, $(U_{n},V_{n})_{n}$ converges in distribution to $(\mathbb{U},\mathbb{V})$ on $\ell^{\infty}(\mathbb{R})\times\ell^{\infty}(0,1)$ by Lemma E.12. Hence, by Slutsky’s lemma, the sequence $(U_{n},V_{n},F_{n})_{n}$ converges in distribution to $(\mathbb{U},\mathbb{V},F)$ on $\ell^{\infty}(\mathbb{R})\times(\ell^{\infty}(0,1))^{2}$. Hence by Lemma E.11 and the continuous mapping theorem we obtain,
@@
Let $U\in\mathcal{C}(\mathbb{R})$ and $(V,F)\in(\ell^{\infty}(0,1))^{2}$. Let $(U_{n},V_{n},F_{n})_{n}$ a sequence in $\ell^{\infty}(\mathbb{R})\times(\ell^{\infty}(0,1))^{2}$ which converges to $(U,V,F)$. Let $K\subset(0,1)$ be a compact set. Let $\varepsilon>0$. The set $\mathopen{}\mathclose{{\left\{F_{n}(x),x\in K,n\in\mathbb{N}}}\right\}$ is bounded since $\mathopen{}\mathclose{{\left\lVert F_{n}-F}}\right\rVert_{\infty,K}\rightarrow 0$ and $(F_{n})_{n}$ and $F$ belongs in $\ell^{\infty}(0,1)$, hence is included in a compact set $K^{\prime}$ of $\mathbb{R}$. Since $U$ is continuous, we have $\eta>0$ such that for all $(x,y)\in{K^{\prime}}^{2}$, $\mathopen{}\mathclose{{\left|x-y}}\right\rvert\leqslant\eta$ implies that $\mathopen{}\mathclose{{\left|U(x)-U(y)}}\right\rvert\leqslant\varepsilon$. Let $N>0$ such that for all $n\geqslant N$, $\mathopen{}\mathclose{{\left\lVert U_{n}-U}}\right\rVert_{\infty,K^{\prime}}\leqslant\varepsilon$ and $\mathopen{}\mathclose{{\left\lVert F_{n}-F}}\right\rVert_{\infty,K}\leqslant\eta$. Then for all $n\geqslant N$,
$\mathopen{}\mathclose{{\left\lVert U_{n}\circ F_{n}-U\circ F}}\right\rVert_{\infty,K}$ $\leqslant\mathopen{}\mathclose{{\left\lVert U_{n}\circ F_{n}-U\circ F_{n}}}\right\rVert_{\infty,K}+\mathopen{}\mathclose{{\left\lVert U\circ F_{n}-U\circ F}}\right\rVert_{\infty,K}$
$\leqslant\mathopen{}\mathclose{{\left\lVert U_{n}-U}}\right\rVert_{\infty,K^{\prime}}+\varepsilon$
@@
where the first $\varepsilon$ appears since for all $x\in K$, $F_{n}(x)\in K^{\prime}$, $F(x)\in K^{\prime}$ and $\mathopen{}\mathclose{{\left|F_{n}(x)-F(x)}}\right\rvert\leqslant\eta$. Furthermore for all $\tilde{K}$ a compact subset of $(0,1)$, $V_{n}-V||_{\infty,\tilde{K}}\rightarrow 0$. Hence, $\Psi$ is continuous at $(U,V,F)$. ∎
Finally, the following result is classical and shows how a convergence in distribution on $D[0,1]$ can imply a uniform convergence (Billingsley, 1999), with the local topology defined in Section 2.3.
Lemma E.12 (Billingsley (1999)).
@@ Appendix F Useful Lemmas
Let $\mathopen{}\mathclose{{\left(s_{1},\cdots,s_{n+1}}}\right)\in{\mathbb{R}\cup\mathopen{}\mathclose{{\left\{+\infty}}\right\}}$, $(w_{k})_{k\in\llbracket n+1\rrbracket}$ be a family of positive weight summing to $1$ and $\mu=\sum_{k\in\llbracket n+1\rrbracket}w_{k}\delta_{s_{k}}$ be a probability measure on $\mathbb{R}\cup\mathopen{}\mathclose{{\left\{+\infty}}\right\$}. Define for all $\alpha\in(0,1)$, $Q_{\alpha}(\mu)=\inf\mathopen{}\mathclose{{\left\{x\in\mathbb{R}:\mu\mathopen{}\mathclose{{\left([-\infty,x]}}\right)\geqslant\alpha}}\right\}$ the $\alpha$-quantile of the probability measure $\mu$. Then, for all $t\in\mathbb{R}$ and $\alpha\in(0,1)$, the two following assertions are equivalent:
@@
$S\coloneqq\max_{k\in\llbracket n+1\rrbracket}\mathopen{}\mathclose{{\left\{s_{k}:s_{k}<t}}\right\}\in\mathbb{R}$